OpenStudy (anonymous):

1,2,4,8,16 what is the formula that would give the value of nth location in sequence? anybody?

OpenStudy (anonymous):

2^(n-1) 2^0 = 1 2^1 = 2 2^2 = 4 If you're using zero indexing (if the first item in the sequence is item 0) then it's just 2^n.

OpenStudy (anonymous):

the first location (that gives the value 1) is location 0. continue from the question.

OpenStudy (anonymous):

how do you get the formula?